Worley Company buys surgical supplies from a variety of manufacturers and then resells and delivers these supplies to hundredRequired: 1. Compute the total revenue that Worley would receive from University and Memorial. 2. Compute the activity rate fComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the acComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the toComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute Worley
Answer #1
University Memorial
Total Revenue      41,730      41,730
2.Activity Rate = Overhead Cost/Total Activity
Customer Deliveries 86
Manual Order Processing 78
Electronic Order Processing 22
Line item picking 1.6
Cost Assigned = Activity Rate*Activity Driver
University Memorial
Customer Deliveries            946        2,236
Manual Order Processing               -          3,744
Electronic Order Processing            286               -  
Line item picking            256            352
Total cost         1,488        6,332
Customer Margin = Revenue - Cost
University Memorial
Customer Margin         1,242       (3,602)

A University =39000*1.07 Memorial =39000*1.07 2 Total Revenue 4 2.Activity Rate = Overhead Cost/Total Activity 5 6 Customer D